# Extending wavelet regularity beyond Gevrey classes

### v1

@@ 3 Wavelet construction
In [7, Theorem 4.4] the authors also constructed a wavelet $\psi\in N_{x}^{2}$. To conclude this paper we compare our regularity with the one proposed by $N^{\sigma}_{x}$.
@@
Let $T(k)=\exp\left\{\rho\,\log^{\frac{\sigma}{\sigma-1}}k\right\},\,k>1,\,\sigma>1,$ be the function that describes the decay rate the decay of $\hat{f}$ in (38). Then by the calculation done in [13, Section 4.1], it follows that $T(k)$ is associated function for $M_{p}=q^{p^{\sigma}}$, for suitable $q>1$ when $1<\sigma\leq 2$.
In particular we can prove the following Lemma.
@@
for suitable $C,\rho^{\prime}>0$ and $|\xi|$ large enough. This proves that $f\in\mathcal{E}_{\sigma}^{\Gamma}$. Inclusion $\mathcal{E}_{\sigma}^{\Gamma}\subset\mathcal{E}_{\sigma}$ follows directly from the Theorem 2.11 b). ∎
@@ 3.1 Illustrations
In this section, we present illustrations of the low-pass filter constructed in the proof of Theorem 1.2, together with the corresponding scaling function and wavelet. All figures were generated in Matlab. The parameters used in (36) and (37) are $\sigma=2$ and $d=\frac{\pi}{12}$.
Figure 1: The plot of the low-pass filter $m_{0}$ given in (37). Note that it features a small bump near $\tfrac{2\pi}{3}$, which is a consequence of controlling its local decay.
@@ 4 Appendix
hence part $b)$ of the Lemma 2.16 and (43) imply that each term of the form (42) tends to 0 when $s\to+\infty$. Finally, this proves $\lim_{x\to 0^{+}}f_{\rho_{\sigma},\sigma}^{(j)}(x)=0$. ∎
@@ 5 Acknowledgments
This research was supported by the Science Fund of the Republic of Serbia, #\#GRANT No. 2727, Global and local analysis of operators and distributions - GOALS. The authors were also supported by the Ministry of Science, Technological Development, and Innovation of the Republic of Serbia - the second and third authors by Grants No. 451-03-137/2025-03/200125 and 451-03-136/2025-03/200125, and the first author by Grant No. 451-03-136/2025-03/200156.

### v2

@@ 3 Wavelet construction
In [8, Theorem 4.4], the authors also constructed a wavelet $\psi\in N_{x}^{2}$ by imposing polynomial estimates on $m_{0}$ in a neighborhood of $\frac{2\pi}{3}$. This approach allows them to obtain improved regularity of the wavelet in the frequency domain. On the other hand, our wavelet has regularity of the class $\mathcal{E}_{\sigma}(\mathbb{R})$ in both the time and frequency domains, as a consequence of controlling $m_{0}$ by the functions given in (3).
To conclude this paper we compare our regularity with the one proposed by $N^{\sigma}_{x}$. We start with the following Remark.
@@
Let $T(k)=\exp\left\{\rho\,\log^{\frac{\sigma}{\sigma-1}}k\right\},\,k>1,\,\sigma>1,$ be the function that describes the decay rate the decay of $\hat{f}$ in (38). Then by the calculation done in [14, Section 4.1], it follows that $T(k)$ is associated function for $M_{p}=q^{p^{\sigma}}$, for suitable $q>1$ when $1<\sigma\leq 2$.
Now we can compare regularities in the following Lemma.
@@
for suitable $C,\rho^{\prime}>0$ and $|\xi|$ large enough. This proves that $f\in\mathcal{E}_{\sigma}^{\Gamma}$. Inclusion $\mathcal{E}_{\sigma}^{\Gamma}\subset\mathcal{E}_{\sigma}$ follows directly from the Theorem 2.11. ∎
@@ 3.1 Illustrations
In this section, we present illustrations of the low-pass filter constructed in the proof of Theorem 3.5, together with the corresponding scaling function and wavelet. All figures were generated in Matlab (the code is available at https://sites.google.com/view/goalsproject/publications). The parameters used in (36) and (37) are $\sigma=2$ and $d=\frac{\pi}{12}$.
Figure 1: The plot of the low-pass filter $m_{0}$ given in (37). Note that it features a small bump near $\tfrac{2\pi}{3}$ (see the smaller piece), that enables control of its local decay.
@@ 4 Appendix
hence Lemma 2.15 $b)$ and (43) imply that each term of the form (42) tends to 0 when $s\to+\infty$. Finally, this proves $\lim_{x\to 0^{+}}f_{\rho_{\sigma},\sigma}^{(j)}(x)=0$. ∎
@@ 5 Acknowledgments
The authors thank N. Teofanov for helpful discussions. This research was supported by the Science Fund of the Republic of Serbia, #\#GRANT No. 2727, Global and local analysis of operators and distributions - GOALS. The authors were also supported by the Ministry of Science, Technological Development, and Innovation of the Republic of Serbia - the second and third authors by Grants No. 451-03-137/2025-03/200125 and 451-03-136/2025-03/200125, and the first author by Grant No. 451-03-136/2025-03/200156.